OpenStudy (anonymous):

Please Help! Has to be finished by tonight. Will fan medal, write a testimonial, anything. Create two radical equations: one that has an extraneous solution, and one that does not have an extraneous solution. Use the equation below as a model.

TheSmartOne (thesmartone):

Where is the equation below?

OpenStudy (anonymous):

one sec

OpenStudy (anonymous):

\[a \sqrt{x+b+c}=d\]

OpenStudy (anonymous):

Use a constant in place of each variable a, b, c, and d. You can use positive and negative constants in your equation.

TheSmartOne (thesmartone):

People have asked this question previously. The links are above...

OpenStudy (anonymous):

Thank you so much

TheSmartOne (thesmartone):

\(\Huge\color{Green}{My~Pleasure}\)

OpenStudy (anonymous):

\[2 \sqrt{x+5+7}=4\] So this is correct? @TheSmartOne

OpenStudy (anonymous):

@TheSmartOne